OpenStudy (anonymous):

Find the geometric mean of 20 and 5. A. 10 B. 4 C. 12.5 D. 25

OpenStudy (anonymous):

would it be C?

OpenStudy (vishweshshrimali5):

Uncle rhakus can u help me too after this question http://openstudy.com/study#/updates/4fab31cbe4b059b524f703e5

OpenStudy (unklerhaukus):

\[G=\sqrt[n]{x_1\times x_2\times \dots x_n}\]

OpenStudy (vishweshshrimali5):

is it A

OpenStudy (unklerhaukus):

\[G=\sqrt[2]{20\times5}=\sqrt{100}=...\neq25\]

OpenStudy (anonymous):

20+5/2=12.5

OpenStudy (vishweshshrimali5):

it is arithmetic mean dear

OpenStudy (unklerhaukus):

@ashwinjohn3 that is the arithmetic mean